# Phase transition on the fluctuation of the structure of random walk ranges

### v1

@@ Abstract.
We investigate fluctuation phenomena for the graph distance and the number of cut points associated with random media arising from the range of a random walk. Our results demonstrate a sequence of dimension-dependent phase transitions in the scaling behavior of these fluctuations, leading to qualitatively different regimes across dimensions lower than six, equal to six, and higher than six. In particular, in six dimensions, we show that the convergence to the normal distribution occurs under a scaling different from that of the standard central limit theorem. We also give the related results for the effective resistance of the random walk range.
@@ 1. Random walk on random walk range
$V({\mathcal{G}}):=\{S_{m}:m\geq 0\},$
$E({\mathcal{G}}):=\{\{S_{m},S_{m+1}\}:m\geq 0\}.$
Originally, Lawler and Burdzy [5] discussed the exponents for the fluctuation of the random walk range and proposed many open problems. Croydon and Shiraishi [11, 20, 21] estimated the graph distance, the effective resistance of $(V({\mathcal{G}}),E({\mathcal{G}}))$ and the number of cut points of the random walk range between $S_{0}$ and $S_{n}$ in $d=4$, and showed that their typical order is $n/(\log n)^{1/2+o(1)}$. In addition, Croydon [7] did similar work in $d\geq 5$, and showed that their typical order is linear. In [22], he discussed the loop erased random walk and other related concepts for the graph of the random walk range in $d=2,3$.
@@
An important tool for analyzing random walks in random media is the effective resistance or graph distance associated with the underlying random media. In particular, the behavior of the effective resistance is closely connected to the spectral dimension of the random subset, which characterizes the long-time decay of return probabilities and provides a key indicator of whether the associated random walk exhibits diffusive or anomalous scaling.
@@
The behavior of graph distance plays a central role in a variety of problems, including volume growth, metric scaling, and universality phenomena in random geometry. In particular, in light of recent developments, it is conjectured that graph distances in several models of random media exhibit non-trivial scaling behavior, potentially related to general universality phenomena. These considerations provide strong motivation for a detailed analysis of graph distances in random media, which is a focus of the present work.
Motivated by these considerations, we aim to understand the fluctuation behavior of the effective resistance and the graph distance in random media. In this paper, we take the range of a random walk as a concrete example of a random medium. According to the results of Croydon [7], in $d\geq 5$ the leading-order behavior is linear regardless of the dimension and does not possess a phase transition, which naturally raises the question of how fluctuations behave around this leading term.
Our analysis reveals that the fluctuation behavior depends sensitively on the dimension. In particular, we find distinct regimes in dimensions four, five, six, and seven and higher. A major phase transition occurs at dimension six: for $d\geq 6$, the fluctuations, when normalized by the square root of the variance, converge in distribution to a Gaussian random variable, whereas in dimensions four and five, the same normalization leads to convergence to zero. We also note that the variance of the normal random variable in $d=6$ that appears in the central limit theorem is non-standard in the sense that it is actually not equal to the variance of the original random variable. Moreover, our proofs suggest that in dimensions four and five, if the fluctuations are rescaled by the appropriate order, they are expected to converge to non-Gaussian random variables.
@@
This corresponds to the energy dissipation of the resistance electrical network on $G$. We also introduce the space
$H^{2}:=\{\,f\in\mathbb{R}^{V}:\mathcal{E}(f,f)<\infty\,\},$
This quantity represents the voltage difference required to send the current from $A$ to $B$ and therefore corresponds to the usual notion of effective resistance in an electrical network. For vertices $x,y\in V$, we write $R_{G}(x,y)$ to denote $R_{G}(\{x\},\{y\})$.
We will now define the notation used to represent the graph distance, the effective resistance, and the number of cut points. Let $X^{1}_{[a,b]}(x,y):=d_{\mathcal{G}_{[a,b]}}(x,y)$, $X^{2}_{[a,b]}(i,j)$ be the number of cut points on $\mathcal{G}_{[a,b]}$ between the time $i$ and $j$, and $X^{3}_{[a,b]}(x,y):=R_{\mathcal{G}_{[a,b]}}(x,y)$. A trivial relationship among these variables is $X^{3}_{[a,b]}(x,y)\leq X^{1}_{[a,b]}(x,y)$ and $X^{2}_{[a,b]}(i,j)\leq X^{3}_{[a,b]}(S_{i},S_{j})$. We sometimes write $X^{i}[a,b]$ as $X^{i}_{[a,b]}(S_{a},S_{b})$ and $X^{i}_{n}=X^{i}_{[0,n]}(S_{0},S_{n})$ for $i=1,3$ and $X^{2}[a,b]$ as $X^{2}_{[a,b]}(a,b)$ and $X^{2}_{n}=X^{2}_{[1,n]}(1,n)$. Since the following proof is the same, we omit the symbol “ii”, wherever it appears. To ease the presentation, we omit hereafter the integer-part symbol $\lceil\cdot\rceil$. If there exists $c\in(0,\infty)$ such that $ca_{n}\leq b_{n}\leq c^{-1}a_{n}$, we write $a_{n}\asymp b_{n}$. In addition, if $a_{n}/b_{n}\to 1$ as $n\to\infty$, we write $a_{n}\sim b_{n}$. If $\log a_{n}/\log b_{n}\to 1$ as $n\to\infty$, we write $a_{n}\approx b_{n}$. $a_{n}\gg b_{n}$ means $a_{n}/b_{n}\to\infty$ as $n\to\infty$ and $a_{n}\ll b_{n}$ means $b_{n}/a_{n}\to\infty$ as $n\to\infty$. We write $f(n)\lesssim g(n)$ if there exists a (deterministic) constant $c>0$ such that $f(n)\leq cg(n)$ for all $n$, and $f(n)\gtrsim g(n)$ if $g(n)\lesssim f(n)$.
First, we will give our main results for the graph distance and the number of cut points:
@@
It holds that, as $n\to\infty$,
$\mathrm{Var}(X_{n})\quad\begin{cases}\sim\alpha^{i}_{d}n&\text{for }d\geq 7,\\ \asymp n\log n&\text{for }d=6,\\ \asymp n^{3/2}&\text{for }d=5,\\ =\frac{n^{2}}{(\log n)^{2\pm o(1)}}&\text{for }d=4,\end{cases}$
for some constant $\alpha^{i}_{d}>0$ for $i=1,2$.
@@
For $d\geq 7$, as $n\to\infty$,
@@
$\frac{X_{n}-\mathbb{E}X_{n}}{\sqrt{\phi(n)}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1).$
@@
$\frac{X_{n}-\mathbb{E}X_{n}}{\sqrt{\mathrm{Var}(X_{n})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}0.$
Next, we will give the result for the effective resistance.
@@
For $d\geq 7$ and $\alpha^{3}_{d}>0$, as $n\to\infty$,
$\mathrm{Var}(X_{n})\sim\alpha^{3}_{d}n.$
@@
$\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n\log n}>0,\quad\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n(\log n)^{2}}<\infty.$
@@
$\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n^{2}/(\log n)^{2+\epsilon}}>0,\quad\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n})}{n^{2}/(\log n)^{1-\epsilon}}<\infty.$
@@
For $d\geq 7$, as $n\to\infty$,
$\frac{X_{n}-\mathbb{E}X_{n}}{\sqrt{\mathrm{Var}(X_{n})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1).$
@@
$\frac{X_{n_{k}}-\mathbb{E}X_{n_{k}}}{\sqrt{\phi(n_{k})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1).$
@@
$\frac{X_{n}-\mathbb{E}X_{n}}{\sqrt{\mathrm{Var}(X_{n})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}0,$
@@
$\frac{X_{n_{k}}-\mathbb{E}X_{n_{k}}}{\sqrt{\mathrm{Var}(X_{n_{k}})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}0.$
Open problem 1.5.
Remark 1.6.
We believe that the stronger results of Propositions 1.1 and 1.2 are true for the effective resistance as well, but we avoid writing the extra technical details for simplicity. Instead, we present a weaker version that holds under weaker assumptions, such as without any correlation assumptions on the cross terms; we believe that the analysis performed in this section is of independent interest and can be readily applicable to any random variable on the random walk range that is governed by heavy tailed behavior.
As a key ingredient of our approach to the analysis of fluctuations, we employ a decomposition of the random walk path. More precisely, we split the trajectory $S[0,n]$ into two subpaths $S[0,n/2]$ and $S[n/2,n]$, and compare the corresponding effective resistances or graph distances. The quantity
X[0,n/2]+X[n/2,n]X[0,n]X[0,n/2]+X[n/2,n]-X[0,n]
naturally arises as error terms.
By iterating this decomposition procedure, we obtain a representation consisting of a sum of finely decomposed random variables, which behave approximately as independent and identically distributed terms, together with accumulated error terms. A key observation is that a phase transition occurs depending on the dimension: while in dimensions $d\geq 6$ the sum of the i.i.d.-type terms dominates the fluctuations, in lower dimensions the error terms become dominant and significantly affect the scaling behavior. Consequently, it is expected that for $d\geq 6$ the fluctuations converge in distribution to a Gaussian random variable, whereas such Gaussian behavior does not persist in lower dimensions.
We outline the main technical difficulties in the proof of our results. For the graph distance and the number of cut points, establishing the validity or failure of the central limit theorem, which constitutes our main theorem, requires a precise and delicate analysis of the variance. In particular, this analysis necessitates a careful control of the correlations among the error terms arising from the decomposition employed in the proof. As a first step, it is necessary to improve the result of [4, Lemma 1.1] by identifying the optimal strategy for events in the large deviation regime that give rise to large error terms. We show that such events correspond to situations in which the random walk forms a large loop. The second step consists in estimating the correlations of these error terms. More precisely, one must show that these correlations are sufficiently small when the time intervals associated with the decomposition that generates the error terms are well separated.
A technical difficulty to show the validity or failure of the central limit theorem is the fact that the large deviation probabilities governing these error terms exhibit heavy-tailed behavior. In fact, a related phenomenon for the decomposition was already observed in earlier work of the authors [1, 2, 12] concerning the capacity of the range of a random walk. In that setting, when analyzing fluctuations via a similar decomposition scheme, a phase transition occurred between dimensions four and five, depending on which term became dominant in the decomposition.
While this qualitative behavior is closely related to the phenomena studied in the present paper, the analysis in our current setting requires substantially different techniques. More specifically, in the previous work the error terms arising from the decomposition became negligible in dimensions $d\geq 5$, whereas in the present case the error terms remain of polynomial order even for $d\geq 6$, necessitating a much more delicate analysis. Moreover, although in dimension four the error terms are leading in both settings, their control in the present work requires global information on the entire path, leading to more intricate estimates.
We also obtained partial results for the effective resistance; however, the resulting bounds are weaker than those for the other quantities. This is due to technical difficulties arising in the estimation of the correlations of the error terms. Nevertheless, even without a detailed analysis of these correlations, once an estimate of the error terms corresponding to an improvement of [4, Lemma 1.1] is available, we are able to obtain an analogy that yields a phase transition for the central limit theorem of the effective resistance under weaker assumptions on the error terms. We believe that this analogy is useful for a broad class of random variables for which a decomposition of the type considered in this work can be carried out.

### v2

@@ Abstract.
We investigate fluctuation phenomena for the graph distance and the number of cut points associated with random media arising from the range of a random walk. Our results demonstrate a sequence of dimension-dependent phase transitions in the scaling behavior of these fluctuations, leading to qualitatively different regimes, with a distinct phase transition in dimension 6. In particular, we remark that convergence in dimension 6 occurs with a non-standard rescaling.
@@ 1. Random walk on random walk range
$V({\mathcal{G}}):=\{S_{m}:m\geq 0\},\quad E({\mathcal{G}}):=\{\{S_{m},S_{m+1}\}:m\geq 0\}.$
Originally, Lawler and Burdzy [5] discussed the exponents for the fluctuations of the random walk range and proposed many open problems. Croydon and Shiraishi [11, 20, 21] estimated the graph distance, the effective resistance of $(V({\mathcal{G}}),E({\mathcal{G}}))$ and the number of cut points of the random walk range between $S_{0}$ and $S_{n}$ in $d=4$, and showed that their typical orders are $n/(\log n)^{1/2+o(1)}$. In addition, Croydon [7] did similar work in $d\geq 5$ and showed that their typical order is linear. In [22], he discussed the loop-erased random walk and other related concepts for the graph of the random walk range in $d=2,3$.
@@
The behavior of graph distance plays a central role in a variety of problems, including volume growth, metric scaling, and universality phenomena in random geometry. In particular, in light of recent developments, it is conjectured that graph distances in several models of random media exhibit non-trivial scaling behavior, potentially related to general universality phenomena. These considerations provide strong motivation for a detailed analysis of graph distances in random media, which is the focus of the present work.
Motivated by these considerations, we aim to understand the fluctuation behavior of the graph distance in random media. In this paper, we take the range of a random walk as a concrete example of a random medium. According to the results of Croydon [7], in $d\geq 5$ the leading-order behavior is linear regardless of the dimension and does not possess a phase transition, which naturally raises the question of how fluctuations behave around this leading term.
Our analysis reveals that the fluctuation behavior depends sensitively on the dimension. In particular, we find distinct regimes in dimensions four, five, six, and seven and higher. A major phase transition occurs at dimension six: for $d\geq 6$, the fluctuations, when normalized by the square root of the variance, converge in distribution to a Gaussian random variable, whereas in dimensions four and five, the same normalization leads to convergence of the fluctuation to zero. We also note that the variance of the normal random variable in $d=6$ that appears in the central limit theorem is non-standard in the sense that it is actually not equal to the variance of the original random variable.
@@
We will now define the notation used to represent the graph distance and the number of cut points. Let $X^{1}_{[a,b]}(x,y):=d_{\mathcal{G}_{[a,b]}}(x,y)$ and $X^{2}_{[a,b]}(i,j)$ be the number of cut points on $\mathcal{G}_{[a,b]}$ between the time $i$ and $j$. A trivial relationship among these variables is $X^{1}_{[a,b]}(S_{i},S_{j})\geq X^{2}_{[a,b]}(i,j)$. We sometimes write $X^{i}[a,b]$ as $X^{i}_{[a,b]}(S_{a},S_{b})$ and $X^{1}_{n}=X^{1}_{[0,n]}(S_{0},S_{n})$ and $X^{2}[a,b]$ as $X^{2}_{[a,b]}(a,b)$ and $X^{2}_{n}=X^{2}_{[1,n]}(1,n)$. Since the following proof is the same, we omit the symbol “ii”, wherever it appears. To ease the presentation, we omit hereafter the integer-part symbol $\lceil\cdot\rceil$. If there exists $c\in(0,\infty)$ such that $ca_{n}\leq b_{n}\leq c^{-1}a_{n}$, we write $a_{n}\asymp b_{n}$. In addition, if $a_{n}/b_{n}\to 1$ as $n\to\infty$, we write $a_{n}\sim b_{n}$. If $\log a_{n}/\log b_{n}\to 1$ as $n\to\infty$, we write $a_{n}\approx b_{n}$. $a_{n}\gg b_{n}$ means $a_{n}/b_{n}\to\infty$ as $n\to\infty$ and $a_{n}\ll b_{n}$ means $b_{n}/a_{n}\to\infty$ as $n\to\infty$. We write $f(n)\lesssim g(n)$ if there exists a (deterministic) constant $c>0$ such that $f(n)\leq cg(n)$ for all $n$, and $f(n)\gtrsim g(n)$ if $g(n)\lesssim f(n)$.
We now present our main results on the fluctuation of the graph distance and the number of cut points.
@@
Consider $i=1,2$. It holds that, as $n\to\infty$,
$\mathrm{Var}(X_{n}^{i})\quad\begin{cases}\sim\alpha^{i}_{d}n&\text{for }d\geq 7,\\ \asymp n\log n&\text{for }d=6,\\ \asymp n^{3/2}&\text{for }d=5,\\ =\frac{n^{2}}{(\log n)^{2\pm o(1)}}&\text{for }d=4,\end{cases}$
for some constant $\alpha^{i}_{d}>0$ with $i=1,2$.
@@
Consider $i=1,2$. For $d\geq 7$, as $n\to\infty$,
$\frac{X_{n}^{i}-\mathbb{E}X_{n}^{i}}{\sqrt{\mathrm{Var}(X_{n}^{i})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1)$
@@
$\frac{X_{n}^{i}-\mathbb{E}X_{n}^{i}}{\sqrt{\phi(n)}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1).$
@@
$\frac{X_{n}^{i}-\mathbb{E}X_{n}^{i}}{\sqrt{\mathrm{Var}(X_{n}^{i})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}0.$
Open problem 1.3.
We emphasize that the theorem above on the fluctuations of the graph distance and the number of cut points demonstrates a key phase transition in dimension 6. While in dimensions strictly greater than 6, one has the standard central limit theorem, this fails in dimensions less than 6. The standard scaling for the fluctuation leads to a trivial random variable in dimensions 4 and 5. In addition, dimension 6 presents a subtle, but deep change in the behavior of the fluctuation; instead of fluctuations on the scale of $\sqrt{n}$, as in higher dimensions, the fluctuation will be of the scale $\sqrt{n\log n}$. Moreover, the precise order of the fluctuation in dimension 6 is not given by the square root of the variance, but differs from the variance by a constant. Indeed, by looking at the proof, one can intuit that the main contribution to the fluctuation in dimensions higher than seven arises primarily from local fluctuations in the random walk range. In dimension 6, the proof demonstrates that the primary source of the fluctuation arises from interactions at intermediate scales of order $\sqrt{n}$. As such, in dimension 6, we observe the first demonstration of the effect of long-range interactions for the graph distance of the random walk range. We remark that the source of this new phase transition is due to the effect of the heavy tailed nature of the long range interactions of the random walk that can cause significant effects on the graph distance.
One of the key ingredients of our proof is to quantify the effect of the long-range interactions of the random walk range by a decomposition. On the highest level, one can measure the effect of the interaction between the first and second halves of the random walk by splitting the trajectory $S[0,n]$ into two subpaths $S[0,n/2]$ and $S[n/2,n]$. Once this is done, one can compare the corresponding graph distances by studying the variable
X[0,n/2]+X[n/2,n]X[0,n],X[0,n/2]+X[n/2,n]-X[0,n],
that naturally arises as an error term.
By iteratively applying this decomposition procedure, we obtain a representation consisting of a sum of finely decomposed random variables that behave approximately as independent and identically distributed terms, together with accumulated error terms. The key question here is to determine which of the terms contributes the most to the variance, and on which scale they operate. In dimensions $d\geq 7$, one can very finely decompose the random variables and show that local contributions will contribute the most to the fluctuations; since the local contributions are independent of each other, this will lead to a Gaussian fluctuation. In dimensions 4 and 5, one will observe that the variance almost entirely comes from the error term at the largest scale; namely, it comes from considering X[0,n/2]+X[n/2,n]X[0,n]X[0,n/2]+X[n/2,n]-X[0,n].
In dimension 6, there becomes a key interplay between the contributions of the error terms and the sum of the independent terms. Indeed, fully decomposing the random walk will not lead to a correct understanding of the fluctuation; rather, the most important contributions to the fluctuation in dimension 6 come from intermediate scales of order $\sqrt{n}$. Due to the heavy tails, the probability that the error term X[0,n/2]+X[n/2]X[0,n]X[0,n/2]+X[n/2]-X[0,n] will make significant contributions to the fluctuation is around $n^{-1}$. At the intermediate scale of size $\sqrt{n}$, we will have $\sqrt{n}$ terms that have a probability $(\sqrt{n})^{-1}$ of contributing non-trivially. Thus, the intermediate scales of size $\sqrt{n}$ are the first scales that start to contribute with high probability to the fluctuation. The total fluctuation of $n\log n$ comes from accumulating all the interactions from scale $\sqrt{n}$ or lower and carefully taking into account their interactions. This is in contrast to the naive computation of the variance, which accumulates some level of contributions from all scales, even those that are not seen in the fluctuation.
Many of the key challenges in determining the fluctuations in the graph distance or the number of cut points require a deep understanding of the nature of the error terms that arise in our analysis. In contrast to the study of many other variables for the random walk, we remark that the number of cut points or the graph distance is a purely discrete quantity. That is, one cannot understand the graph distance by considering the scaling limit of the random walk to the Brownian motion, and then analyzing the appropriate statistics of the Brownian motion. As a result, there is a lack of closed formulas that can be used to analyze the error terms.
Another aspect of the analysis of the error terms is that they are very heavy tailed. This occurs because large values of this error term are governed by events that, while rather unlikely, involve very long-range dependencies in the random walk. This behavior of the error terms causes two fundamental problems for our analysis. First, long-range dependencies cause the error terms at different scales to become correlated. Secondly, one now has to take into account a balance between the rarity of a given interaction and the effect of a given error term to determine whether it contributes to the scale of our fluctuations.
Solving these problems requires a very precise analysis of the covariance structure of the error terms. In particular, many previous estimates we found in the literature were insufficient for our purposes. In higher dimensions, we sharpen a result of [4, Lemma 1.1] to identify the optimal order of fluctuation of the given error terms and remove an important $O(\log n)$ factor. A delicate analysis on the correlation of cut times and long-range intersections of the random walk allows us to determine the covariance structures of the cross terms in dimensions $d\geq 5$ (observe the proofs of Lemma 4.1 and Lemma 5.1). In dimension 4, we used a link between intersection properties and the capacity of the random walk, combined with a multiscale analysis in a series of lemmas, in order to decouple the interaction between two subpaths of a random walk and the graph distance on either subpath and obtain optimal estimates on the variance of the cross terms in Section 6. Finally, in the proof of Theorem 1.2, we performed the careful balancing act that allowed us to exactly determine the correct order of fluctuation; this involved showing that, due to the heavy tails, the highest scales occur too rarely, and then carefully analyzing the correlations between the cross terms at intermediate scales. Our proof reveals that the intermediate scales of order $\sqrt{n}$ determined the main contribution to the fluctuation in six dimensions.
In addition to our analysis of the graph distance, we present, for independent interest, an analysis of quantities for the random walk range via our decomposition strategy, whose error terms exhibit the same heavy tailed behavior as the graph distance or the number of cut points. In this more general analysis, we do not make any assumptions about the correlations of the cross term. As a result, it will not be possible to obtain the same sharp results as one could obtain for the graph distance; however, our analysis demonstrates that heavy tailed behavior, by itself, is enough to demonstrate a type of phase transition similar to those shown for the graph distance. Indeed, we believe that these more general results would be helpful for analyzing other random variables of interest, and, furthermore, demonstrate the key role of the size of the heavy tails in the cross terms in creating a phase transition for the fluctuations.
@@
Let $X^{3}_{[a,b]}$ be any collection positive random variables that can be written as a measurable function, $\mathcal{F}(S[a,b])$, with respect to the random walk range $S[a,b]$ (with $X^{3}_{n}:=X^{3}_{[0,n]}$). Here, we treat $S[a,b]$ as a subgraph of $\mathbb{Z}^{d}$. We further assume that $\mathcal{F}$ is translation invariant and can be bounded from above by the volume. Define $E_{n}:=X^{3}_{n}+X^{3}_{[n,2n]}-X^{3}_{[0,2n]}$ and assume further that $E_{n}$ satisfies the bounds
$\mathbb{P}(E_{n}\geq l)\asymp l^{-(d/2-2)},\quad d\geq 5,$
and $E_{n}$ is always positive, that is, $\mathcal{F}$ is subadditive. For $d=4$, the right-hand side above does not meaningfully characterize heavy tailed behavior, so the natural analogs of the heavy tailed assumption are those found in Assumption 6.16.
Then, for $d\geq 7$ and $\alpha^{3}_{d}>0$, as $n\to\infty$,
$\mathrm{Var}(X_{n}^{3})\sim\alpha^{3}_{d}n.$
@@
$\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n}^{3})}{n\log n}>0,\quad\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n}^{3})}{n(\log n)^{2}}<\infty.$
@@
$\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n}^{3})}{n^{2}/(\log n)^{2+\epsilon}}>0,\quad\limsup_{n\to\infty}\frac{\mathrm{Var}(X_{n}^{3})}{n^{2}/(\log n)^{1-\epsilon}}<\infty.$
@@
Let $X_{n}^{3}$ satisfy the assumptions outlined in the previous lemma. For $d\geq 7$, as $n\to\infty$,
$\frac{X_{n}^{3}-\mathbb{E}X_{n}^{3}}{\sqrt{\mathrm{Var}(X_{n}^{3})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1)$
@@
$\frac{X_{n_{k}}^{3}-\mathbb{E}X_{n_{k}}^{3}}{\sqrt{\phi(n_{k})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,1).$
@@
$\frac{X_{n}^{3}-\mathbb{E}X_{n}^{3}}{\sqrt{\mathrm{Var}(X_{n}^{3})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}0,$
@@
$\frac{X_{n_{k}}^{3}-\mathbb{E}X_{n_{k}}^{3}}{\sqrt{\mathrm{Var}(X_{n_{k}}^{3})}}\stackrel{{\scriptstyle d}}{{\rightarrow}}0.$